# Online Detecting LLM-Generated Texts via Sequential Hypothesis Testing by Betting

### v2

@@ 3 Our Algorithm
We have covered most of the underlying algorithmic design principles of our online method for detecting LLMs, and we are now ready to introduce our algorithm, which is shown in Algorithm 1. Compared to existing works on sequential hypothesis testing via betting (e.g., Shekhar and Ramdas (2023); Chugg et al. (2023)), which assume knowledge of a bound on the magnitude of the “coin outcome" $g_{t}$ in the learner’s wealth dynamic (1) for all time steps before the testing begins (e.g., assuming for all $t$, $\phi(x_{t})$ and $\phi(y_{t})$ are always in $[0,1]$), we relax this assumption, which is motivated by the concern that a good estimate of the bound for the coin outcome might not be available before the sequential testing starts. Specifically, we consider the scenario where an upper bound on $|g_{t+1}|$ at round $t+1$, which is denoted by $d_{t+1}$, is available before updating $\theta_{t+1}$ at each round $t$. Our algorithm then plays a point in the decision space $\mathcal{K}_{t+1}$ that guarantees the learner’s wealth remains a non-negative supermartingale (Step 15 in Algorithm 1). We note that if the bound of the output of the underlying score function $\phi(\cdot)$ is known a priori, this scenario holds naturally. Otherwise, we can estimate an upper bound for $|g_{t}|$ for all $t$ based on the first few time steps and execute the algorithm thereafter. One approach is to set the estimate as a conservatively large constant, e.g., twice the maximum value observed in the first few time steps. We observe that this estimate works for our algorithm with most of the score functions $\phi(\cdot)$ that we consider in the experiments. On the other hand, we note that a tighter bound $d_{t}$ will lead to a faster time to reject $H_{0}$ when the unknown source is an LLM, as indicated by the following propositions.
@@
$\mathbb{E}[\tau]\lesssim\frac{d_{*}^{3}}{\Delta^{2}}\cdot\log\left(\frac{d_{*}^{(3+\frac{1}{\gamma})}}{\Delta^{2}\alpha}\right),$ (2)
where $\Delta:=|\mu_{x}-\mu_{y}|$, $d_{*}:=\max\limits_{t\geq 1}\lvert d_{t}\rvert$ with $d_{t}\geq|g_{t}|$, and $\gamma$ satisfies $\gamma\leq\frac{1}{2}\min\{\frac{d_{t}}{G_{t}},1\}$ with $G_{t}:=\max_{\theta\in\mathcal{K}_{t}}|\nabla\ell_{t}(\theta)|$ denoting the upper bound of the gradient $\nabla\ell_{t}(\theta)$.
@@
Consequently, the dynamic of the wealth evolves as $W_{t}^{A}=W_{t-1}^{A}\left(1-\theta_{t}(g_{t}-\epsilon)\right)$ and $W_{t}^{B}=W_{t-1}^{B}\left(1-\theta_{t}(-g_{t}-\epsilon)\right)$ respectively, where $g_{t}=\phi({x}_{t})-\phi({y}_{t})$. We note that both $g_{t}-\epsilon$ and $-g_{t}-\epsilon$ are within the interval $[-d_{t}-\epsilon,d_{t}-\epsilon]$. The composite hypothesis is motivated by the fact that, in practice, even if both sequences of texts $x_{t}$ and $y_{t}$ are human-written, they may have been written by different individuals. Therefore, it might be more reasonable to allow for a small difference $\epsilon$ in their means when defining the null hypothesis $H_{0}$.
@@
$\mathbb{E}[\tau]\lesssim\frac{(d_{*}+\epsilon)^{3}}{(\Delta-\epsilon)^{2}}\cdot\log\left(\frac{(d_{*}+\epsilon)^{(3+\frac{1}{\gamma})}}{(\Delta-\epsilon)^{2}\alpha}\right).$ (3)
Remark 2. Proposition 2 indicates that even if there is a difference $\epsilon$ in mean scores between texts written by different humans, the probability that the source is incorrectly declared by Algorithm 3 as an LLM can be controlled below $\alpha$. Furthermore, the bound on $\mathbb{E}[\tau]$ implies that smaller $\epsilon$ and larger $\Delta$ will result in a shorter time to reject $H_{0}$.
@@ 4 Experiments
Parameters of Our Algorithm. All the experiments in the following consider the setting of the composite hypothesis. For the step size $\gamma$, we simply follow the related works (Cutkosky and Orabona, 2018; Chugg et al., 2023; Shekhar and Ramdas, 2023) and let $1/\gamma=2/(2-\log 3)$. We consider two scenarios of sequential hypothesis testing in the experiments. The first scenario (oracle) assumes that one has prior knowledge of $d_{*}$ and $\epsilon$, and the performance of our algorithm in this case could be considered as an ideal outcome that it can achieve. For simulating this ideal scenario in the experiments, we let $\epsilon$ be the absolute difference between the mean scores of XSum texts and 2024 Olympic news, which are datasets of human-written texts. The second scenario considers that we do not have such knowledge a priori, and hence we have to estimate $d_{*}$ (or $d_{t}$) and specify the value of $\epsilon$ using the samples collected in the first few times steps, and then the hypothesis testing is started thereafter. In our experiments, we use the first $10$ samples from each sequence of $x_{t}$ and $y_{t}$ and set $d_{t}$ to be a constant, which is twice the value of $\max_{s\leq 10}\lvert\phi(x_{s})-\phi(y_{s})\rvert$. For estimating $\epsilon$, we obtain scores for $20$ texts sampled from the XSum dataset and randomly divided them into two groups, and set $\epsilon$ to twice the average absolute difference between the empirical means of these two groups across $1000$ random shuffles.
Experimental Results. The experiments evaluate the performance of our method and baselines under both $H_{0}$ and $H_{1}$. As there is inherent randomness from the observed samples of the texts in the online setting, we repeat $1000$ runs and report the average results over these $1000$ runs. Specifically, we report the false positive rate (FPR) under $H_{0}$, which is the number of times the source of $y_{t}$ is incorrectly declared as an LLM when it is actually human, divided by the total of $1000$ runs. We also report the average time to reject the null under $H_{1}$ (denoted as Rejection Time $\tau$), which is the time our algorithm takes to reject $H_{0}$ and correctly identify the source when $y_{t}$ is indeed generated by an LLM. More precisely, the rejection time $\tau$ is the average time at which either $W_{t}^{A}$ or $W_{t}^{B}$ exceeds $2/\alpha$ before $T$; otherwise, $\tau$ is set to $T=500$, regardless of whether it rejects $H_{0}$ at $T$, since the the time budget runs out. The parameter value $d_{t}\leftarrow d_{*}$ in Scenario 1 (oracle) is shown in Table 2, and the value for $\epsilon$ can be found in Table 1 in the appendix. For the estimated $\epsilon$ and $d_{t}$ of each sequential testing in Scenario 2, they are displayed in Table 5 and Table 6 respectively in the appendix. Our method and the baselines require specifying the significance level parameter $\alpha$. In our experiments, we try $20$ evenly spaced values of the significance level parameter $\alpha$ that ranges from $0.005$ to $0.1$ and report the performance of each one.
@@
Figure 3: Averaged results of Scenario 2, where our algorithm has to use the first $10$ samples to specify $d_{t}$ and $\epsilon$ before starting the algorithm. The plots are about the average of the rejection time under $H_{1}$ (i.e., the average time to detect LLMs) and the false positive rate under $H_{0}$ for $10$ different score functions and $20$ different values of the significance level parameter $\alpha$ when using two different scoring models, Neo-2.7 (a) and Gemma-2B (b).
@@
Figure 3 shows the empirical results of our algorithm under Scenario 2, where it has to use the first few samples to specify $d_{t}$ and $\epsilon$ before starting the algorithm. Under this scenario, our algorithm equipped with most of the score functions still performs effectively. We observe that our algorithm with 1) Fast-DetectGPT as the score function $\phi(\cdot)$ and Neo-2.7 as the language model for computing the score, and with 2) Likelihood as the score function $\phi(\cdot)$ and Gemma-2B for computing the value of $\phi(\cdot)$ have the best performance under this scenario. Compared to the first case where the oracle of $d_{*}$ and $\epsilon$ is available and exploited, these two configurations only result in a slight degradation of the performance under Scenario 2, and we note that our algorithm can only start updating after the first $10$ time steps under this scenario. We observe that the bound of $d_{*}$ that we estimated using the samples collected from first $10$ time steps is significantly larger than the tightest bound of $d_{*}$ in most of the runs where we refer the reader to Table 2, 6 in the appendix for details, which explains why most of the configurations under Scenario 2 need a longer time to detect LLMs, as predicted by our propositions. We also observe that the configurations with two supervised classifiers (RoBERTa-based and RoBERTa-large) and the combinations of a couple of score functions and the scoring model Gemma-2B do not strictly control FPRs across all significance levels. This is because the estimated $d_{t}$ for these score functions is not large enough to ensure that the wealth $W_{t}$ remains nonnegative at all time points $t$. That is, we observed $2d_{t}<\lvert\phi(x_{t})-\phi(y_{t})\rvert+\epsilon$ for some $t$ in the experiments, and hence the wealth $W_{t}$ is no longer a non-negative supermartingale, which prevents the application of Ville’s inequality to guarantee a level-$\alpha$ test. Nevertheless, our algorithm with eight score functions that utilize the scoring model Neo-2.7 can still effectively control type-I error and detect LLMs by around $t=300$.
@@
In Appendix H, we provide more experimental results, including those using existing datasets from Bao et al. (2023) for simulating the sequential testing, where our algorithm on these datasets also performs effectively. Moreover, we found that the rejection time is influenced by the relative magnitude of $\Delta-\epsilon$ and $d_{t}-\epsilon$, as predicted by our propositions, and the details are provided in Appendix G. From the experimental results, when the knowledge of $d_{*}$ and $\epsilon$ is not available beforehand, as long as the estimated $d_{*}$ and $\epsilon$ guarantee a nonnegative supermartingale, and the estimated $\epsilon$ is greater than or equal to the actual absolute difference in the empirical mean scores of two sequences of human texts, our algorithm can maintain a sequential valid level-$\alpha$ test and efficiently detect LLMs.
@@
Comparisons with Baselines. In this part, we use the score function of Fast-DetectGPT and scoring model Neo-2.7 to get text scores, and then compare the performance of our method with two baselines that adapt the fixed-time permutation test to the sequential hypothesis setting. Batch sizes $k\in\{25,50,100,250,500\}$ are considered for the baselines. We set the estimated $\epsilon$ and $d_{*}$ values the same as in Scenario 2. The baselines are also implemented in a manner to conduct the composite hypothesis test. We observe a significant difference between the scores of XSum texts and machine-generated news, which causes the baselines of the permutation test to reject the null hypothesis most of the time immediately after receiving the first batch. This in turn results in the nearly vertical lines observed in the left subfigure of Figure 4(a) and Figure 4(b), where the averaged rejection time across 1000 repeated tests closely approximates the batch size $k$ for each of the 20 significance levels $\alpha$. On the other hand, we observe that when $H_{0}$ is true, the baselines might not be a valid $\alpha$-test, even with a corrected significance level. This arises from the increased variability of text scores introduced by smaller batch sizes, which results in observed absolute differences in means that may exceed the $\epsilon$ value under $H_{0}$. Our method can quickly detect an LLM while keeping the false positive rates (FPRs) below the specified significance levels, which is a delicate balance that can be difficult for the baselines to achieve. Without prior knowledge of the $\epsilon$ value, the baselines of permutation tests may fail to control the type-I error with small batch sizes and cannot quickly reject the null hypothesis while ensuring that FPRs remain below $\alpha$, unlike our method.
@@ 5 Conclusion
In this paper, we demonstrate that our algorithm, which builds on the score functions of offline detectors, can rapidly determine whether a stream of text is generated by an LLM and provides strong statistical guarantees. Specifically, it controls the type-I error rate below the significance level, ensures that the source of LLM-generated texts can eventually be identified, and guarantees an upper bound on the expected time to correctly declare the unknown source as an LLM under a mild assumption. Although the choice of detector can influence the algorithm’s performance and some parameters related to text scores need to be predefined before receiving the text, our experimental results show that most existing detectors provide effective score functions, and our method performs well in most cases when using estimated values of parameters based on text scores from the first few time steps. To further enhance its efficacy, it may be worthwhile to design score functions tailored to the sequential setting, improve parameter estimations with scores from more time steps, and study the trade-offs between delaying the start of testing and obtaining more reliable estimates. Moreover, our algorithm could potentially be used as an effective tool to promptly identify and mitigate the misuse of LLMs in generating texts, such as monitoring social media accounts that disseminate harmful information generated by LLMs, rapidly detecting sources of fake news generated by LLMs on public websites, and identifying users who post LLM-generated comments to manipulate public opinion. Exploring these applications could be a promising direction.
Acknowledgments

### v3

@@ 3 Our Algorithm
We have covered most of the underlying algorithmic design principles of our online method for detecting LLMs, and we are now ready to introduce our algorithm, which is shown in Algorithm 1. Compared to existing works on sequential hypothesis testing via betting (e.g., Shekhar and Ramdas (2023); Chugg et al. (2023)), which assume knowledge of a bound on the magnitude of the “coin outcome" $g_{t}$ in the learner’s wealth dynamic (1) for all time steps before the testing begins (e.g., assuming for all $t$, $\phi(x_{t})$ and $\phi(y_{t})$ are always in $[0,1]$), we relax this assumption, which is motivated by the concern that a good estimate of the bound for the coin outcome might not be available before the sequential testing starts. Specifically, we consider the scenario where an upper bound on $|g_{t+1}|$ at round $t+1$, which is denoted by $d_{t+1}$, is available before updating $\theta_{t+1}$ at each round $t$. Our algorithm then plays a point in the decision space $\mathcal{K}_{t+1}$ that guarantees the learner’s wealth remains a non-negative supermartingale (Step 12 in Algorithm 1). We note that if the bound of the output of the underlying score function $\phi(\cdot)$ is known a priori, this scenario holds naturally. Otherwise, we can estimate an upper bound for $|g_{t}|$ for all $t$ based on the first few time steps and execute the algorithm thereafter. One approach is to set the estimate as a conservatively large constant, e.g., twice the maximum value observed in the first few time steps. We observe that this estimate works for our algorithm with most of the score functions $\phi(\cdot)$ that we consider in the experiments. On the other hand, we note that a tighter bound $d_{t}$ will lead to a faster time to reject $H_{0}$ when the unknown source is an LLM, as indicated by the following propositions, where we note that the analysis of the test power requires the assumption that the data are i.i.d.
@@
$\mathbb{E}[\tau]=\mathcal{O}\left(\frac{d_{*}^{2}}{\Delta^{2}}\ln\left(\frac{d_{*}}{\alpha\Delta}\right)+\frac{d_{*}^{4}}{\Delta^{4}}\right),$
where $\Delta:=|\mu_{x}-\mu_{y}|$, $d_{*}:=\max\limits_{t\geq 1}d_{t}$ with $d_{t}\geq|g_{t}|$, and $\gamma=\frac{1}{2}\min\{\frac{d_{t}}{G_{t}},1\}$ with $G_{t}:=\max_{\theta\in\mathcal{K}_{t}}|\nabla\ell_{t}(\theta)|$ denoting the upper bound of the gradient $\nabla\ell_{t}(\theta)$.
@@
Consequently, the dynamic of the wealth evolves as $W_{t}^{A}=W_{t-1}^{A}\cdot\left(1-\theta_{t}(g_{t}-\epsilon)\right)$ and $W_{t}^{B}=W_{t-1}^{B}\cdot\left(1-\theta_{t}(-g_{t}-\epsilon)\right)$ respectively, where $g_{t}=\phi({x}_{t})-\phi({y}_{t})$. We note that both $g_{t}-\epsilon$ and $-g_{t}-\epsilon$ are within the interval $[-d_{t}-\epsilon,d_{t}-\epsilon]$. The composite hypothesis is motivated by the fact that, in practice, even if both sequences of texts $x_{t}$ and $y_{t}$ are human-written, they may have been written by different individuals. Therefore, it might be more reasonable to allow for a small difference $\epsilon>0$ in their means when defining the null hypothesis $H_{0}$.
@@
$\mathbb{E}[\tau]=\mathcal{O}\left(\frac{(d_{*}+\epsilon)^{2}}{(\Delta-\epsilon)^{2}}\ln\left(\frac{d_{*}+\epsilon}{\alpha(\Delta-\epsilon)}\right)+\frac{(d_{*}+\epsilon)^{4}}{(\Delta-\epsilon)^{4}}\right),$
where $\Delta:=|\mu_{x}-\mu_{y}|$, $d_{*}:=\max\limits_{t\geq 1}d_{t}$ with $d_{t}\geq|g_{t}|$, $\gamma=\frac{1}{2}\min\{\frac{2d_{t}}{G_{t}},1\}$ with $G_{t}:=\max_{\theta\in\mathcal{K}_{t}}|\nabla\ell_{t}(\theta)|$ denoting the upper bound of the gradient $\nabla\ell_{t}(\theta)$, and $\epsilon$ is a nonnegative constant satisfying $\epsilon\leq d_{t}$ for all $t$.
Remark 2. We note that when the alternative hypothesis $H_{1}$ is true, we have $\Delta>\epsilon$. Proposition 2 indicates that even if there is a difference $\epsilon$ in mean scores between texts written by different humans, the probability that the source is incorrectly declared by Algorithm 3 as an LLM can be controlled below $\alpha$. Furthermore, the bound on $\mathbb{E}[\tau]$ implies that smaller $\epsilon$ and larger $\Delta$ will result in a shorter time to reject $H_{0}$.
@@ 4 Experiments
Parameters of Our Algorithm. All the experiments in the following consider the setting of the composite hypothesis. For the step size $\gamma$, we simply follow the related works (Cutkosky and Orabona, 2018; Chugg et al., 2023; Shekhar and Ramdas, 2023) and let $1/\gamma=2/(2-\ln 3)$. We consider two scenarios of sequential hypothesis testing in the experiments. The first scenario (oracle) assumes that one has prior knowledge of $d_{t}$ (or $d_{*}$) and $\epsilon$, and the performance of our algorithm in this case could be considered as an ideal outcome that it can achieve. For simulating this ideal scenario in the experiments, we let $\epsilon$ be the absolute difference between the mean scores of XSum texts and 2024 Olympic news, which are datasets of human-written texts. The second scenario considers that we do not have such knowledge a priori, and hence we have to estimate $d_{t}$ (or $d_{*}$) and specify the value of $\epsilon$ using the samples collected in the first few times steps, and then the hypothesis testing is started thereafter. In our experiments, we use the first $10$ samples from each sequence of $x_{t}$ and $y_{t}$ and set $d_{t}$ to be a constant, which is twice the value of $\max_{s\leq 10}\lvert\phi(x_{s})-\phi(y_{s})\rvert$. For estimating $\epsilon$, we obtain scores for $20$ texts sampled from the XSum dataset and randomly divide them into two groups, and set $\epsilon$ to twice the average absolute difference between the empirical means of these two groups across $1000$ random shuffles.
Experimental Results. The experiments evaluate the performance of our method and baselines under both $H_{0}$ and $H_{1}$. As there is inherent randomness from the observed samples of the texts in the online setting, we repeat $1000$ runs and report the average results over these $1000$ runs. Specifically, we report the false positive rate (FPR) under $H_{0}$, which is the number of times the source of $y_{t}$ is incorrectly declared as an LLM when it is actually human, divided by the total of $1000$ runs. We also report the average time to reject the null under $H_{1}$ (denoted as Rejection Time $\tau$), which is the time our algorithm takes to reject $H_{0}$ and correctly identify the source when $y_{t}$ is indeed generated by an LLM. More precisely, the rejection time $\tau$ is the average time at which either $W_{t}^{A}$ or $W_{t}^{B}$ exceeds $2/\alpha$ before $T$; otherwise, $\tau$ is set to $T=500$, regardless of whether it rejects $H_{0}$ at $T$, since the the time budget runs out. The parameter value $d_{t}$ in Scenario 1 (oracle) is shown in Table 2, and the value for $\epsilon$ can be found in Table 1 in the appendix. For the estimated $\epsilon$ and $d_{t}$ of each sequential testing in Scenario 2, they are displayed in Table 5 and Table 6 respectively in the appendix. Our method and the baselines require specifying the significance level parameter $\alpha$. In our experiments, we try $20$ evenly spaced values of the significance level parameter $\alpha$ that ranges from $0.005$ to $0.1$ and report the performance of each one.
@@
Figure 3: Averaged results of Scenario 2, where our algorithm has to use the first few samples to specify $d_{t}$ and $\epsilon$ before starting the algorithm. The plots are about the average of the rejection time under $H_{1}$ (i.e., the average time to detect LLMs) and the false positive rate under $H_{0}$ for $10$ different score functions and $20$ different values of the significance level parameter $\alpha$ when using two different scoring models, Neo-2.7 (a) and Gemma-2B (b).
@@
Figure 3 shows the empirical results of our algorithm under Scenario 2, where it has to use the first few samples to specify $d_{t}$ and $\epsilon$ before starting the algorithm. Under this scenario, our algorithm equipped with most of the score functions still performs effectively. We observe that our algorithm with 1) Fast-DetectGPT as the score function $\phi(\cdot)$ and Neo-2.7 as the language model for computing the score, and with 2) Likelihood as the score function $\phi(\cdot)$ and Gemma-2B for computing the value of $\phi(\cdot)$ have the best performance under this scenario. Compared to the first case where the oracle of $d_{t}$ and $\epsilon$ is available and exploited, these two configurations only result in a slight degradation of the performance under Scenario 2, and we note that our algorithm can only start updating after the first $10$ time steps under this scenario. We observe that the bound of $d_{t}$ that we estimated using the samples collected from first $10$ time steps is significantly larger than the tightest bound of $d_{t}$ in most of the runs where we refer the reader to Table 2, 6 in the appendix for details, which explains why most of the configurations under Scenario 2 need a longer time to detect LLMs, as predicted by our propositions. We also observe that the configurations with two supervised classifiers (RoBERTa-based and RoBERTa-large) and the combinations of a couple of score functions and the scoring model Gemma-2B do not strictly control FPRs across all significance levels. This is because the estimated $d_{t}$ for these score functions is not large enough to ensure that the wealth $W_{t}$ remains nonnegative at all time points $t$. That is, we observed $2d_{t}<\lvert\phi(x_{t})-\phi(y_{t})\rvert+\epsilon$ for some $t$ in the experiments, and hence the wealth $W_{t}$ is no longer a non-negative supermartingale, which prevents the application of Ville’s inequality to guarantee a level-$\alpha$ test. Nevertheless, our algorithm, when using most score functions that utilize the scoring model Neo-2.7, can still effectively control type-I error and detect LLMs around $t=300$.
@@
In Appendix H, we provide more experimental results, including those using existing datasets from Bao et al. (2023) for simulating the sequential testing, where our algorithm on these datasets also performs effectively. Moreover, we found that the rejection time is influenced by the relative magnitude of $\Delta-\epsilon$ and $d_{t}-\epsilon$, as predicted by our propositions, and the details are provided in Appendix G. From the experimental results, when the knowledge of $d_{t}$ and $\epsilon$ is not available beforehand, as long as the estimated $d_{t}$ and $\epsilon$ guarantee a nonnegative supermartingale, and the estimated $\epsilon$ lies between the actual mean score difference of human-written texts and that of human-written vs LLM-generated texts, our algorithm can maintain a sequential valid level-$\alpha$ test and efficiently detect LLMs.
Comparisons with Baselines. In this part, we use the score function of Fast-DetectGPT and scoring model Neo-2.7 to get text scores, and then compare the performance of our method with two baselines that adapt the fixed-time permutation test to the sequential hypothesis setting. Batch sizes $k\in\{25,50,100,250,500\}$ are considered for the baselines. We set the estimated $\epsilon$ and $d_{t}$ values the same as in Scenario 2. The baselines are also implemented in a manner to conduct the composite hypothesis test.
@@
We observe a significant difference between the scores of XSum texts and machine-generated news, which causes the baselines of the permutation test to reject the null hypothesis most of the time immediately after receiving the first batch. This in turn results in the nearly vertical lines observed in the left subfigure of Figure 4(a) and Figure 4(b), where the averaged rejection time across 1000 repeated tests closely approximates the batch size $k$ for each of the 20 significance levels $\alpha$. On the other hand, we observe that when $H_{0}$ is true, the baselines might not be a valid $\alpha$-test, even with a corrected significance level. This arises from the increased variability of text scores introduced by smaller batch sizes, which results in observed absolute differences in means that may exceed the $\epsilon$ value under $H_{0}$. Our method can quickly detect an LLM while keeping the false positive rates (FPRs) below the specified significance levels, which is a delicate balance that can be difficult for the baselines to achieve. Without prior knowledge of the $\epsilon$ value, the baselines of permutation tests may fail to control the type-I error with small batch sizes and cannot quickly reject the null hypothesis while ensuring that FPRs remain below $\alpha$, unlike our method.
@@ 5 Conclusion
In this paper, we demonstrate that our algorithm, which builds on the score functions of offline detectors, can rapidly determine whether a stream of text is generated by an LLM and provides strong statistical guarantees. Specifically, it controls the type-I error rate below the significance level, ensures that the source of LLM-generated texts can eventually be identified, and guarantees an upper bound on the expected time to correctly declare the unknown source as an LLM under a mild assumption. Although the choice of detector can influence the algorithm’s performance and some parameters related to text scores need to be predefined before receiving the text, our experimental results show that most of the existing detectors we use provide effective score functions, and our method performs well in most cases when using estimated values of parameters based on text scores from the first few time steps. To further enhance its efficacy, it may be worthwhile to design score functions tailored to the sequential setting, improve parameter estimations with scores from more time steps, and study the trade-offs between delaying the start of testing and obtaining more reliable estimates. Moreover, our algorithm could potentially be used as an effective tool to promptly identify and mitigate the misuse of LLMs for text generation, such as monitoring social media accounts that disseminate harmful information generated by LLMs, rapidly detecting sources of LLM-generated fake news on public websites, and identifying users who post machine-generated comments to manipulate public opinion. Exploring these applications could be a promising direction.
Finally, we note that a recent work of Chen and Wang (2025) introduces optimistic interior-point methods that enable updates over the full interior of the decision space. Incorporating such advanced betting strategies into the testing-by-betting framework presents a promising future direction for improving online detection of machine-generated texts.
Acknowledgements